OpenStudy (anonymous):

simplify 2.325 X 10^6/9.3 X 10^9

OpenStudy (anonymous):

1/4000

OpenStudy (anonymous):

(2.325*10^6)/(9.3*10^9)=(2.325/9.3)*10^-3 =2.5*10^-4

OpenStudy (anonymous):

because...... this equals 2.325*1000000/9.3*1000000000 =(2.325/9.3=1/4),(10^6/10^9=1/10^3)so...=1/4*1/10^3=1*10/4*10^3=1/4000

OpenStudy (anonymous):

Rationalize the fractions:\[\frac{93}{40} * 10^6/\frac{93}{10} *10^9=250*10^{12} \]